eliminate the fractions in the following equation:
x(3/4) + y(2/3) + z(5/6) = 2

Answer by mananth(16946)

x(3/4) + y(2/3) + z(5/6) = 2
LCM of the left side = 12
x(9)+y(8)+z(10) / 12 = 2
x(9)+y(8)+z(10)= 12 * 2
9x+8y+10z=24